OpenStudy (anonymous):

@TheSmartOne please Sally has only nickels and dimes in her money box. She knows that she has less than $20 in the box. Let x represent the number of nickels in the box and y represent the number of dimes in the box. Which of the following statements best describes the steps to graph the solution to the inequality in x and y? Draw a dashed line to represent the graph of 5x + 10y = 2000, and shade the portion above the line for positive values of x and y. Draw a dashed line to represent the graph of 10x + 5y = 2000, and shade the portion above the line for positive values of x and y.

OpenStudy (anonymous):

I think it is C

OpenStudy (anonymous):

@TheSmartOne please help

OpenStudy (anonymous):

@sleepyjess please help

myininaya (myininaya):

x is number of nickels and we know nickels are .05 dollars (also called 5 cents-whatever) y is number of dimes and we know dimes are .10 dollars (also called 10 cents-whatever) So anyways I'm going to use dollars here so I can have the same units as the 20 dollars on the other side the equation. So we know .05x+.10y=20

myininaya (myininaya):

you can get rid of the decimals by multiplying both sides by 100

myininaya (myininaya):

resulting in one the equations given above

OpenStudy (anonymous):

so it is 5x+10y=20

myininaya (myininaya):

not exactly

myininaya (myininaya):

you multiply one side by 100

myininaya (myininaya):

but not the other

OpenStudy (anonymous):

so its 5x+.10=20

myininaya (myininaya):

That is actually getting a little worse than before.

OpenStudy (anonymous):

im lost

myininaya (myininaya):

You multiplied the left hand side by 100 but you did not multiply the 20 by 100

OpenStudy (anonymous):

oh so it is 5+.10=2000

myininaya (myininaya):

Your equation that you first typed was almost correct All I said was that you needed to multiply the other side by 100 as well But now you are giving an equation that doesn't have x and y and for some reason you still have .1 instead of 10

myininaya (myininaya):

We had initially .05x+.10y=20

myininaya (myininaya):

multiply both sides by 100

OpenStudy (anonymous):

5x+10y=2000

myininaya (myininaya):

100(.05x)+100(.10y)=100(20)

myininaya (myininaya):

yes!

OpenStudy (anonymous):

and that is it

myininaya (myininaya):

now you are actually given that 5x+10y will amount to something less than 2000

myininaya (myininaya):

correct?

OpenStudy (anonymous):

true but can it be equal to also

myininaya (myininaya):

so instead of a equal sign by a less than sign

myininaya (myininaya):

well it doesn't say less than or equal to 2000

myininaya (myininaya):

it says less than

myininaya (myininaya):

5x+10y<2000

myininaya (myininaya):

we could actually simplify this equation more but your choices look like they do not do that

OpenStudy (anonymous):

oh ok but all of the answer choices show = signs

myininaya (myininaya):

yeah they are telling you how to graph 5x+10y<2000

myininaya (myininaya):

we need the less than sign to know which part of the graph to shade and to know if we have a broken line or a solid line

OpenStudy (anonymous):

it is broken and below the line right

OpenStudy (anonymous):

Sally has only nickels and dimes in her money box. She knows that she has less than $20 in the box. Let x represent the number of nickels in the box and y represent the number of dimes in the box. Which of the following statements best describes the steps to graph the solution to the inequality in x and y? Draw a dashed line to represent the graph of 5x + 10y = 2000, and shade the portion above the line for positive values of x and y. Draw a dashed line to represent the graph of 10x + 5y = 2000, and shade the portion above the line for positive values of x and y. Draw a dashed line to represent the graph of 5x + 10y = 2000, and shade the portion below the line for positive values of x and y. Draw a dashed line to represent the graph of 10x – 5y = 2000, and shade the portion below the line for positive values of x and y.

myininaya (myininaya):

if there was actually an equal sign we would have a solid line but we have no equal sign in 5x+10y<2000 so we definitely know the line is broken

myininaya (myininaya):

yep

myininaya (myininaya):

so I think you said choice C at the very get go

OpenStudy (anonymous):

so the answer would be the third choice or C

myininaya (myininaya):

and that would be correct

OpenStudy (anonymous):

thankyou

myininaya (myininaya):

np